OpenStudy (anonymous):

what are the solutions of the equations x^2+14x=-130? (1point) a.2,-16 b.10,-13 c.-2,16 d.no solution

OpenStudy (anonymous):

i think its d but im not sure

OpenStudy (anonymous):

You can just plug in the points and if you get left side = right side that's a solution.

OpenStudy (anonymous):

b i guess this is super hard am i right

pooja195 (pooja195):

is this from the quiz?

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

no @poorja195 its not im trieg to see how i got it wrong

OpenStudy (anonymous):

wow are u guys just ignoring me ok fine i wont explain

OpenStudy (anonymous):

i sent you a messege @pflorido

OpenStudy (anonymous):

with a tiny bit of practice you can see that there is no solution to this right away half of 14 is 7 and \(7^2=49\) so we can go right to \[(x+7)^2=-130+49\]

OpenStudy (anonymous):

the left side is a perfect square, the right side is negative therefore there is no real solution because any square must be greater than or equal to zero

OpenStudy (anonymous):

that's what i thought but wasn't sur bc i took this quiz and it said i was wrong